An 85-kg runner is accelerating at a rate of 2 m/s^2 and a 65-kg runner at a rate of 3 m/s^2. If the heavier runner started at a speed of 1 m/s while the lighter runner started off stationary, then how long is it before the runners have the same momentum?


1
Expert's answer
2022-01-02T18:28:49-0500

Answer

Mass of first runner m1=85Kg

Initial velocity u1=1"\frac{m}{sec}"

Accereation a1=2 "\frac{m}{sec^2}"

For second runner

Mass m2=65Kg

Initial velocity u2=0(staionary)

Acceleration a2="3\frac{m}{sec}"

For given case momentum is same after t time velocities

For first runner

"v_1=u_1+at\\=1+2t"

Similiarly

"v_2=u_2+at\\=0+3t"

Now given is that momentum is same after t time.

"m_1v_1=m_2v_2\\"

"\frac{v_1}{v_2}=\frac{m_2}{m_1}=\frac{65}{85}"


Putting velocity value

"\frac{3t}{1+2t}=\frac{13}{17}"

Now solving this equation we get value of t=3.4 sec.